OpenStudy (emilylizabeth):

Calculus- derivatives. I have two questions i need help with. My teacher told me i got a part of each answer wrong. Can someone please tell me which part and what the answer is. First one: -22x^-5 - 17x^-11/ 21x^-4 = 22x^-2/21 + 17x^-8/3. Second one: 7x^2+5x^9/4x^7= 5x/2- 35/4x^6

jimthompson5910 (jim_thompson5910):

can you post a screenshot of the full problem?

OpenStudy (emilylizabeth):

@jim_thompson5910 Yes i can.

OpenStudy (emilylizabeth):

Its questions 2 and 3. I got the rest by myself.

jimthompson5910 (jim_thompson5910):

Ok let's focus on problem 2 for now

OpenStudy (emilylizabeth):

Im sorry! it was 3 and 4!

jimthompson5910 (jim_thompson5910):

ok so let's focus on problem 3

jimthompson5910 (jim_thompson5910):

I'm assuming problem 3 is supposed to look like this \[\Large \frac{-22x^{-5}-17x^{-11}}{21x^{-4}}\]

jimthompson5910 (jim_thompson5910):

or is it supposed to be this? \[\Large -22x^{-5}-\frac{17x^{-11}}{21x^{-4}}\]

OpenStudy (emilylizabeth):

Its the first one.

jimthompson5910 (jim_thompson5910):

before we do any derivatives, let's simplify the expression

jimthompson5910 (jim_thompson5910):

We can break up the fraction to go from this \[\Large \frac{-22x^{-5}-17x^{-11}}{21x^{-4}}\] to this \[\Large \frac{-22x^{-5}}{21x^{-4}}-\frac{17x^{-11}}{21x^{-4}}\] agreed?

OpenStudy (emilylizabeth):

Yes.

jimthompson5910 (jim_thompson5910):

how would you simplify each fraction? any ideas?

OpenStudy (emilylizabeth):

I feel like it would be the fifth rule?

jimthompson5910 (jim_thompson5910):

the second actually

jimthompson5910 (jim_thompson5910):

we're dividing power expressions of the same base (x) so we subtract exponents

OpenStudy (emilylizabeth):

So -4 minus -4?

jimthompson5910 (jim_thompson5910):